# Universal Analog Computation: Fraïssé limits of dynamical systems

### v1

@@ 10. Algebroidal categories of deterministic systems
An obvious strategy is to choose a ‘natural’ countable class $S$ of shifts and consider the full subcategory $\mathsf{DS}^{\mathsf{ef}}(S)$ of the category $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ of deterministic systems that is ‘generated’ by $S$. More precisely, a deterministic system $(X,T)$ is generated by $S$ if $(X,T)$ is the inverse limit of a countable inverse sequence of shifts in $S$. Recall remark 4.6 that a colimit along ef-pairs is the same thing as a limit along factors. So our formal definition is this:
Definition 10.1.
Intuitively, the shifts in $S$ are our basic building blocks with which we can build more complex systems, namely the $\omega$-proshifts of over $S$.
@@
Such constructions of categories are common in mathematics, because of their desirable properties. For example, in topology, Stone spaces (aka profinite spaces) are the limits of finite discrete spaces; similarly for profinite groups and profinite graphs. In domain theory, bifinite domains are the colimits along embedding-projection pairs of finite partial orders with a least element. The general idea is that of an Ind-completion (resp., Pro-completion), i.e., the ‘closure’ of a category under filtered colimits (resp., limits) [59, ch. 6]. Given remark 4.6, we can use Ind- and Pro-completions interchangeably here. However, one difference in our definition is that we only consider $\omega$-chains and not any filtered diagram. We stress this by adding the prefix $\omega$’ to ‘proshift’.
In this section, we show that the sketched strategy indeed yields algebroidal subcategories of deterministic systems:
@@
For the proof, we need the Curtis-Lyndon-Hedlund Theorem (see, e.g., [71, thm. 6.2.9]) in the following version. To fix some terminology: Let $X$ be a shift over alphabet $A$. Write $B_{n}(X)$ for the set of all words $w\in A^{n}$ that occur in some $x\in X$ (i.e., there is $k$ such that $w=(x(k),\ldots,x(k+n-1))$). For a word $w\in A^{n}$, we call $C_{w}(X):=\{x\in X:w=(x(0),\ldots,x(n-1))\}$ the cylinder set of $w$ in $X$.
@@
Finally, we need a basic category-theoretic fact: that $\mathsf{DS}^{\mathsf{ef}}(S)$ is closed under taking colimits of $\omega$-chains. In other words, a colimit of systems which are colimits of shifts is itself a colimit of shifts. That is well-known for Ind-completions [59, ch. 6]. A sketch of an elementary proof is in appendix A.6.
@@
11. Proshifts of finite type and sofic proshifts
Theorem 10.3 provides a general strategy to build algebroidal categories of deterministic systems: first, choose a countable class of shits, and then close it under (co-) limits along $\omega$-chains. The most studied classes of shifts are (1) the class $\mathsf{SFT}$ of shifts of finite type and (2) the class $\mathsf{SOF}$ of sofic shifts [71, ch. 2 and 3, respectively].
The shifts of finite type are those shifts $X$ described by a finite list of forbidden words, i.e., $X$ is the set of those sequences $x$ over the alphabet $A$ of $X$ that do not contain a word $w$ from a finite set $F$ of words over $A$. Equivalently, shifts of finite type are those shifts whose sequences arise as the sequences of edges of infinite paths in a finite graph. The sofic shifts are those shifts that are factors of shifts of finite type. Equivalently, sofic shifts are those shifts whose sequences arise as the sequences of labels of infinite paths in a finite graph whose edges are labeled. This makes sofic shifts analogous to regular languages in automata theory.
In this section, we apply our strategy to those classes of shifts, so we consider the following categories of deterministic systems.
@@ 11. Proshifts of finite type and sofic proshifts
Both $\mathsf{SFT}$ and $\mathsf{SOF}$ are countable (as they are characterized by finite graphs). Below, we show that they also have the joint embedding property 1 and the amalgamation property 2 mentioned in theorem 10.3. As a consequence, we get our second main universality result, which also adds a perhaps surprising isomorphism.
Corollary 11.2.
There is a unique universal and homogeneous $\omega$-proshift of finite type, and there is a unique universal and homogeneous sofic $\omega$-proshift. Moreover, these two systems are isomorphic. So, for brevity, we call it the universal proshift.
The proof is in subsection 11.1. Then subsection 11.2 investigates the universal proshift and avenues for future work.
11.1. The universal proshift
In this subsection, we prove a sequence of propositions which imply corollary 11.2. The main lemma is the following.
@@ 11.1. The universal proshift
Write $Z:=Y_{0}\times_{f_{0},f_{1}}Y_{1}$. For $w_{0}\in B_{0}^{n}$ and $w_{1}\in B_{1}^{n}$, we also write $\langle w_{0},w_{1}\rangle$ for the word in $(B_{0}\times B_{1})^{n}$ whose $k$-th element is $(w_{0}(k),w_{1}(k))$. For $i\in\{0,1\}$, let $F_{i}$ be a set of $B_{i}$-words such that $Y_{i}$ is the set of $B_{i}$-sequences that avoid $F_{i}$. [^14] It is a basic fact that shifts can equivalently be defined via a set $F$ of avoided words [71, p. 5 and 27]. (The special case where $F$ is finite yields the shifts of finite type.) Define:
@@
As noted, the first part of corollary 11.2 now follows from this proposition together with theorem 10.3. We now establish the second part about the isomorphism.
@@ 11.2. The nature of the universal proshift
Another important concept in dynamical systems theory is the shadowing property [80]. Intuitively, it says that when we numerically simulate an orbit starting in $x_{0}$ and computing, at each step $x_{n}$, the next $x_{n+1}$ with some error from the true $T(x_{n})$, then there is a true orbit close to the simulated orbit. Clearly, such a stability is again welcome from a computational perspective. Formally, the property is defined on a compact metric space $X$ with metric $d$ and continuous function $T:X\to X$. (But it can also be defined purely topologically.) A sequence $(x_{i})$ in $X$ is a $\delta$-pseudo-orbit if $d(x_{i+1},T(x_{i}))<\delta$ for all $i\in\omega$. A point $x\in X$ $\epsilon$-shadows a sequence $(x_{i})$ in $X$ if $d(x_{i},T(x))<\epsilon$ for all $i\in\omega$. Then $(X,T)$ has the shadowing property if, for all $\epsilon>0$, there is $\delta>0$ such that every $\delta$-pseudo-orbit is $\epsilon$-shadowed by some point.
@@
Future work. However, there is much more to be investigated in the categories $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ and $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$. So we now sketch four avenues for future work.
@@
Second, the latter question about genericity is particularly interesting (see again [64, 27, 28, 50, 51, 69]). It also connects to questions about computability, as [51] shows: There are effective actions of $\mathbb{Z}^{1}$ on Cantor space with dense conjugacy class. However, for dimensions $d\geq 2$, there are no dense effective actions of $\mathbb{Z}^{d}$ on Cantor space, while there are non-effective ones. Other computability-theoretic aspects of symbolic dynamics are studied in [36], and in [6] also inverse limits of shifts of finite type occur. In this vein, one might also ask for a concrete description of the universal profshift, not just an existence results (cf. [2]).
@@
Fourth, given the close connection between sofic shifts and automata, it is natural to investigate if there is a similarly fruitful connection between our sofic $\omega$-proshifts and profinite automata [84]. Here duality-theoretic tools might also be useful (see [38] and [39, ch. 8]): they consider the Boolean algebra of regular languages (i.e., those languages that are recognized by some automaton) over a given alphabet and with certain residuation operations, and they show that dual to this Boolean algebra is the free profinite monoid. For further connections between profinite semigroups and shift spaces, see [4].
@@ 12. Conclusion
Moreover, it is pertinent to clarifying the connection between our existence result on universal analog computation (which includes ‘learning systems’ like neural networks) to results like the No-Free-Lunch theorems stating the non-existence of a universal learner.
Finally, at a philosophical level, one should clarify the implication of universality and genericity results on the question—which we had to leave open—whether a dynamical system counts as an analog computer. We have mentioned the density and genericity results about actions on the Cantor space, which are within a longer tradition in ergodic theory on understanding which properties a ‘typical’ dynamical system has (see, e.g., [43] or [50] for a short history). Beyond deterministic systems and closer to nondeterministic systems, [35] probabilistically constructed a countable graph that, with probability 1, yields Fraïssé’s universal homogenous graph; and [33] provides a related probabilistic construction for universal domains in domain theory. So one may try to make precise the question whether being an analog computer (whatever this means exactly) is a typical property. Does the fact that the generic automorphism on Cantor space is effective have any bearing on this question? What about the fact that this fails for higher dimensions? Note that algorithmic randomness also studies such questions of ‘typicality’ [31].
@@ A.1. Proofs from section 4
Closed-valued: For $x\in X$, $\varphi(x)$ is closed, so the image $\psi[\varphi(x)]$ is closed by the corollary to the closed graph theorem. Upper-hemicontinuous: The composition of upper-hemicontinuous multifunctions is again upper-hemicontinuous [3, thm. 17.23]. If $F$ and $G$ are total, then, for $x\in X$, $\varphi(x)$ is nonempty, so $\psi\circ\varphi(x)=\psi[\varphi(x)]\neq\emptyset$. ∎
@@ A.4. Proofs from section 7
A.5. Proofs from section 9
@@
A.6. Proofs from section 10

### v2

@@ 9. Algebroidal categories of deterministic systems
As described in section 4.1, our strategy is to choose a countable class $S$ of shifts and consider the full subcategory $\mathsf{DS}^{\mathsf{ef}}(S)$ of $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ consisting of those systems that are limits of shifts in $S$. We defined $\mathsf{DS}^{\mathsf{ef}}(S)$ in definition 4.3 and called its objects $\omega$-proshifts over $S$. Thus, intuitively, the shifts in $S$ are our basic building blocks with which we can build more complex systems, namely the $\omega$-proshifts of over $S$.
@@
Such constructions of categories are common in mathematics, because of their desirable properties. For example, in topology, Stone spaces (aka profinite spaces) are the limits of finite discrete spaces; similarly for profinite groups and profinite graphs. In domain theory, bifinite domains are the colimits along embedding-projection pairs of finite partial orders with a least element. The general idea is that of an Ind-completion (resp., Pro-completion), i.e., the ‘closure’ of a category under filtered colimits (resp., limits) [68, ch. 6]. Given remark 3.13, we can use Ind- and Pro-completions interchangeably here. However, one difference in our definition is that we only consider $\omega$-chains and not any filtered diagram. We stress this by adding the prefix $\omega$’ to ‘proshift’.
In this section, we show that this strategy indeed yields algebroidal subcategories of deterministic systems:
@@
For the proof, we need the Curtis-Lyndon-Hedlund Theorem (see, e.g., [81, thm. 6.2.9]) in the following version. To fix some terminology: Let $X$ be a shift over alphabet $A$. Write $B_{n}(X)$ for the set of all words $w\in A^{n}$ that occur in some $x\in X$ (i.e., there is $k$ such that $w=(x(k),\ldots,x(k+n-1))$). For a word $w\in A^{n}$, we call $C_{w}(X):=\{x\in X:w=(x(0),\ldots,x(n-1))\}$ the cylinder set of $w$ in $X$.
@@
Finally, we need a basic category-theoretic fact: that $\mathsf{DS}^{\mathsf{ef}}(S)$ is closed under taking colimits of $\omega$-chains. In other words, a colimit of systems which are colimits of shifts is itself a colimit of shifts. That is well-known for Ind-completions [68, ch. 6]. A sketch of an elementary proof is in appendix A.5.
@@
10. The universal proshift
Theorem 9.2 provides a general strategy to build algebroidal categories of deterministic systems: first, choose a countable class of shifts, and then close it under (co-) limits along $\omega$-chains. As mentioned in example 3.2, the two most studied classes of shifts are the class $\mathsf{SFT}$ of shifts of finite type and the class $\mathsf{SOF}$ of sofic shifts [81, ch. 2 and 3, respectively]. Indeed, both $\mathsf{SFT}$ and $\mathsf{SOF}$ are countable (as they are characterized by finite graphs). In this section, we apply our strategy to those classes of shifts and obtain our second universality result, namely theorem 4.2 stating the existence of the universal proshift, i.e., the deterministic system that is the universal and homogeneous system both in $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ and in $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$. The proof is in subsection 10.1. Then subsection 10.2 investigates the universal proshift and avenues for future work.
10.1. Proof of theorem 4.4
First, let us give a name to $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ and $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$.
@@ 10. The universal proshift
From theorem 9.2, we know that $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ and $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$ are algebroidal. The next lemma is the key to show the joint embedding property and the amalgamation property.
@@
Write $Z:=Y_{0}\times_{f_{0},f_{1}}Y_{1}$. For $w_{0}\in B_{0}^{n}$ and $w_{1}\in B_{1}^{n}$, we also write $\langle w_{0},w_{1}\rangle$ for the word in $(B_{0}\times B_{1})^{n}$ whose $k$-th element is $(w_{0}(k),w_{1}(k))$. For $i\in\{0,1\}$, let $F_{i}$ be a set of $B_{i}$-words such that $Y_{i}$ is the set of $B_{i}$-sequences that avoid $F_{i}$. [^21] It is a basic fact that shifts can equivalently be defined via a set $F$ of avoided words [81, p. 5 and 27]. (The special case where $F$ is finite yields the shifts of finite type.) Define:
@@
Hence both $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ and $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$ have a universal and homogeneous object. To establish theorem 4.2 it remains to show that these two objects are isomorphic:
@@ 10.2. The nature of the universal proshift
Another important concept in dynamical systems theory is the shadowing property [91]. Intuitively, it says that when we numerically simulate an orbit starting in $x_{0}$ and computing, at each step $x_{n}$, the next $x_{n+1}$ with some error from the true $T(x_{n})$, then there is a true orbit close to the simulated orbit. Clearly, such a stability is again welcome from a computational perspective. Formally, the property is defined on a compact metric space $X$ with metric $d$ and continuous function $T:X\to X$. (But it can also be defined purely topologically.) A sequence $(x_{i})$ in $X$ is a $\delta$-pseudo-orbit if $d(x_{i+1},T(x_{i}))<\delta$ for all $i\in\omega$. A point $x\in X$ $\epsilon$-shadows a sequence $(x_{i})$ in $X$ if $d(x_{i},T^{i}(x))<\epsilon$ for all $i\in\omega$. Then $(X,T)$ has the shadowing property if, for all $\epsilon>0$, there is $\delta>0$ such that every $\delta$-pseudo-orbit is $\epsilon$-shadowed by some point.
@@
Future work. There is much more to be investigated in the categories $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ and $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$. So we now sketch four avenues for future work.
@@
Second, the latter question about genericity is particularly interesting (see again [72, 35, 34, 55, 57, 79]). It also connects to questions about computability, as [57] shows: There are effective actions of $\mathbb{Z}^{1}$ on Cantor space with dense conjugacy class. However, for dimensions $d\geq 2$, there are no dense effective actions of $\mathbb{Z}^{d}$ on Cantor space, while there are non-effective ones. Other computability-theoretic aspects of symbolic dynamics are studied in [43], and in [9] also inverse limits of shifts of finite type occur. In this vein, one might also ask for a concrete description of the universal proshift, not just an existence results (cf. [3]).
@@
Fourth, given the close connection between sofic shifts and automata, it is natural to investigate if there is a similarly fruitful connection between our sofic $\omega$-proshifts and profinite automata [95]. Here duality-theoretic tools might also be useful (see [46] and [45, ch. 8]): they consider the Boolean algebra of regular languages (i.e., those languages that are recognized by some automaton) over a given alphabet and with certain residuation operations, and they show that dual to this Boolean algebra is the free profinite monoid. For further connections between profinite semigroups and shift spaces, see [6].
@@ 11. Conclusion
Moreover, it is pertinent to clarifying the connection between our existence result on universal analog computation (which includes ‘learning systems’ like neural networks) to results like the No-Free-Lunch theorems stating the non-existence of a universal learner (see [111] for a study of universal prediction methods).
Finally, at a philosophical level, one should clarify the implication of universality and genericity results on the question—which we had to leave open—whether a dynamical system counts as an analog computer. We have mentioned the density and genericity results about actions on the Cantor space, which are within a longer tradition in ergodic theory on understanding which properties a ‘typical’ dynamical system has (see, e.g., [50] or [55] for a short history). Beyond deterministic systems and closer to nondeterministic systems, [42] probabilistically constructed a countable graph that, with probability 1, yields Fraïssé’s universal homogenous graph; [40] provides a related probabilistic construction for universal domains in domain theory; and [47, sec. 3.1] provide a concrete model of the projectively universal profinite graph. So one may try to make precise the question whether being an analog computer (whatever this means exactly) is a typical property. Does the fact that the generic automorphism on Cantor space is effective have any bearing on this question? What about the fact that this fails for higher dimensions? Note that algorithmic randomness also studies such questions of ‘typicality’ [38].
@@ A.1. Proofs from section 3
Remark A.1 (Zero-dimensionalizing and compactifying).
Starting with a ‘Polish system’ $(X,T)$, i.e., where $X$ is a Polish space and $T:X\to X$ is a continuous function, we want to construct a—in a sense best possible—system $(Y,S)$, with $Y$ a compact and zero-dimensional Polish space and $S:Y\to Y$ continuous, together with an injective and equivariant function $\eta:X\to Y$ that is continuous with respect to a finer Polish topology on $X$.
We use a result from [27] on how to zero-dimensionalize and compactify Polish spaces. To state this precisely, let us introduce their terminology: A based Polish space is a pair $(X,\mathcal{B})$ where $X$ is a Polish space and $\mathcal{B}$ is a countable base for the topology of $X$. A base-preserving map $f$ from $(X,\mathcal{B})$ to $(Y,\mathcal{C})$ is a function $f:X\to Y$ such that $f^{-1}(C)\in\mathcal{B}$ for all $C\in\mathcal{C}$ (so $f$ is continuous). A compact zero-dimensional Polish space $X$ can canonically be regarded as a based Polish space by taking the collection $\mathrm{Clp}(X)$ of all clopen subsets of $X$ as the base. Next, [27] consider the following categories:
  • $\mathsf{Pol}_{\mathsf{cz}}$: compact zero-dimensional Polish spaces with continuous maps.
  • $\mathsf{Pol}^{\flat}_{z}$: based Polish spaces $(X,\mathcal{B})$, where $X$ is zero-dimensional and $\mathcal{B}$ is a Boolean algebra, with base-preserving maps.
  • $\mathsf{Pol}^{\flat}$: based Polish spaces with base-preserving maps.
Finally, [27] show that (1) $\mathsf{Pol}_{\mathsf{cz}}$ is a reflective subcategory of $\mathsf{Pol}^{\flat}_{z}$ with reflector $c$ and (2) $\mathsf{Pol}^{\flat}_{z}$, in turn, is a coreflective subcategory of $\mathsf{Pol}^{\flat}$ with coreflector $z$. For (1), given a based space $(X,\mathcal{B})$ in $\mathsf{Pol}^{\flat}_{z}$, the compactification $c(X,\mathcal{B})$ is the compact zero-dimensional space $Y$ obtained as the limit of the diagram of finite spaces $P$, where $P$ is a finite partition of $X$ with partition cells in $\mathcal{B}$. The unit $\eta_{(X,\mathcal{B})}:(X,\mathcal{B})\to Y$ is the canonical embedding of $X$ into its compactification $Y$. For (2), given a based space $(X,\mathcal{B})$ in $\mathsf{Pol}^{\flat}$, the zero-dimensionalization $z(X,\mathcal{B})$ is the based space $(X,\overline{\mathcal{B}})$ on the same underlying set $X$ but whose topology is generated by the base $\overline{\mathcal{B}}$, which is the Boolean algebra generated by $\mathcal{B}$. The counit $\epsilon_{(X,\mathcal{B})}:z(X,\mathcal{B})\to(X,\mathcal{B})$ simply is the identity function on $X$.
Now, we use this result to turn a Polish system $(X,T)$ into a deterministic system in our sense:
  • Step 0. Pick a countable base $\mathcal{B}$ of $X$. For example, if $X=\mathbb{R}^{n}$, pick the open balls centered on rational vectors with rational radii. Next close $\mathcal{B}$ under the dynamics, i.e., consider $\mathcal{B}_{T}:=\{T^{-k}(B):B\in\mathcal{B},k\in\omega\}$. Then $(X,\mathcal{B}_{T})$ is a based Polish space. (This step is unnecessary if one thinks that Polish systems should make their choice of dynamically closed base explicit.)
  • Step 1. Take the zero-dimensionalization $z(X,\mathcal{B}_{T})=(X,\overline{\mathcal{B}_{T}})$. Note that, by construction, $T:z(X,\mathcal{B}_{T})\to(X,\mathcal{B}_{T})$ is base-preserving. So, by the adjunction, it lifts to a dynamics $T:z(X,\mathcal{B}_{T})\to z(X,\mathcal{B}_{T})$.
  • Step 2. Compactify $z(X,\mathcal{B}_{T})$ into $Y:=c(z(X,\mathcal{B}_{T}))$, with the embedding $\eta_{X}:=\eta_{z(X,\mathcal{B}_{T})}:X\to Y$. Consider the map $\eta_{X}\circ T:X\to Y$, i.e., viewing the result of the dynamics $T$ in the larger, compactified space $Y$. By the adjunction, there is a unique commuting map $c(T)$:
    $z(X,\mathcal{B}_{T})$$c(z(X,\mathcal{B}_{T}))$$z(X,\mathcal{B}_{T})$$c(z(X,\mathcal{B}_{T}))$$\scriptstyle{T}$$\scriptstyle{\eta_{X}}$$\scriptstyle{c(T)}$$\scriptstyle{\eta_{X}}$
    So $S:=c(T):Y\to Y$ extends the dynamics $T$ on $X$.
In sum, we have the larger system $(Y,S)$ and an injective function $\eta_{X}:X\to Y$, which is equivariant by the above diagram and continuous with respect to the finer Polish topology on $X$ generated by $\overline{\mathcal{B}_{T}}$. The fact that $c$ and $z$ are adjoint functors provides a sense in which this $(Y,S)$ is optimal. [^25] Similar ideas are used in [61, sec. 5.3].
@@
Closed-valued: For $x\in X$, $\varphi(x)$ is closed, so the image $\psi[\varphi(x)]$ is closed by the corollary to the closed graph theorem. Upper-hemicontinuous: The composition of upper-hemicontinuous multifunctions is again upper-hemicontinuous [5, thm. 17.23]. If $F$ and $G$ are total, then, for $x\in X$, $\varphi(x)$ is nonempty, so $\psi\circ\varphi(x)=\psi[\varphi(x)]\neq\emptyset$. ∎
@@ A.3. Proofs from section 6
Proof of lemma 6.10.
A.4. Proofs from section 8
@@
A.5. Proofs from section 9